OpenStudy (anonymous):

Last one! Find the infinite sum, if it exists. THANK YOU! (LOOK AT RESPONSE)

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\left(\begin{matrix}3 \\ 4\end{matrix}\right)^{n}\]

OpenStudy (anonymous):

Same formula as before .. now you tell me what is a and r!